# LayeredMAPF: a decomposition of MAPF instance without compromising solvability

### v1

@@ VI-D Push and Swap
Hierarchical Cooperative A* (HCA*) [silver2005cooperative] is a suboptimal and incomplete prioritized MAPF algorithm that decouples into a series of single-agent searches. The individual searches are performed in three-dimensional space-time, taking into account the planned routes of other agents. Essentially, HCA* decomposes the MAPF instance into subproblems that involve only one agent. HCA* also calculates a distance table for each agent to determine their priority.
@@
On average, layered HCA* has a slightly higher time cost than raw HCA* ($1.3\times 10^{4}$ ms $>$ $1.1\times 10^{4}$ ms) and a similar success rate (layered HCA*: 0.68, raw HCA*: 0.64) as HCA* solves each agent separately, and decomposition does not improve the efficiency of HCA*. However, layered HCA* requires less memory space than raw HCA* ($2.9\times 10^{2}$ MB $<$ $5.4\times 10^{2}$ MB), as layered HCA* only stores the distance table of the current subproblem’s agents, while raw HCA* needs to store the distance table of all agents.
Regarding path quality, layered HCA* has a slightly larger makespan ($2.4\times 10^{2}$ $>$ $2.1\times 10^{2}$) and SOC ($4.6\times 10^{4}$ $>$ $3.8\times 10^{4}$). Because HCA* solves agents in order of decreasing distance to the target, raw HCA* solves agents with longer paths first, limiting the makespan of these agents’ solutions, while layered HCA* breaks this order as agents with long paths may not be in the same subproblem and may not be solved first.
@@
VI-F Priority Inheritance with Backtracking$+$
Priority Inheritance with Backtracking$+$ (PIBT$+$) [okumura2022priority] is an incomplete and suboptimal MAPF algorithm. Its high level assigns a unique priority to each agent at every timestep, while its low level moves each agent one step at every timestep. Consequently, all intermediate states are saved until all agents reach their targets. PIBT$+$ avoids the occurrence of the same state of agents twice in the solution. And the code [^8] https://github.com/Kei18/pibt2 for PIBT$+$ provides no interfaces to avoid conflicts with external paths, making it a parallel MAPF method.
In general, layered PIBT$+$ takes more time than raw PIBT$+$ ($2.0\times 10^{3}$ ms $>$ $6.9\times 10^{2}$ ms). As mentioned in Section V, the decomposition of MAPF instances with 800 to 1000 agents may cost 1s to 3s, and these instances play a major role in contributing to the mean time cost of the MAPF method. Although decomposition reduces the time cost of PIBT$+$, the time cost of decomposition offsets the reduced time cost and thus increases the overall time cost.
Layered PIBT$+$ has a similar success rate compared to raw PIBT$+$ (layered PIBT$+$: 0.99, raw PIBT$+$: 0.99), as both layered PIBT$+$ and raw PIBT$+$ typically take less than 30s in most cases. In terms of memory usage, layered PIBT$+$ requires less memory space than raw PIBT$+$ (6.8 MB $<$ 18 MB), as decomposition reduces the number of intermediate states by reducing the number of agents solved simultaneously.
Regarding solution quality, as PIBT$+$ is a parallel MAPF method, layered PIBT$+$ introduces extra wait actions, resulting in solutions with higher makespan ($9.5\times 10^{3}$ $>$ $4.3\times 10^{2}$) and SOC ($3.1\times 10^{6}$ $>$ $1.2\times 10^{5}$) compared to raw PIBT$+$’s solutions.
@@ VI-F Priority Inheritance with Backtracking$+$
VI-G Lazy Constraints Addition for MAPF
Lazy Constraints Addition for MAPF (LaCAM) [okumura2023lacam] is a complete and suboptimal method. More details about LaCAM can be found in Section II. LaCAM avoids the occurrence of the same state of agents twice in the solution. The code [^9] https://github.com/Kei18/lacam for LaCAM provides no interfaces to avoid conflicts with external paths, making it a parallel MAPF method.
In general, layered LaCAM has slightly lower time cost than raw LaCAM ($3.6\times 10^{3}$ ms $<$ $3.8\times 10^{2}$ ms). Similar to PIBT$+$, the decomposition of MAPF instances with 800 to 1000 agents may cost 1s to 3s, and these instances play a major role in determining the time cost of the MAPF method. Although decomposition reduces the time cost of LaCAM, the time cost of decomposition offsets the reduced time cost, resulting in no significant difference in the overall time cost.
Layered LaCAM has a slightly better success rate compared to raw LaCAM (layered LaCAM: 0.93, raw LaCAM: 0.88), as both layered LaCAM and raw LaCAM typically take less than 30s in most cases. In terms of memory usage, layered LaCAM requires less memory space than raw LaCAM (12 MB $<$ 18 MB), as decomposition reduces the number of intermediate states by reducing the number of agents solved simultaneously.
Regarding solution quality, layered LaCAM introduces extra wait actions, resulting in solutions with higher makespan ($9.3\times 10^{3}$ $>$ $3.2\times 10^{2}$) and SOC ($3.1\times 10^{6}$ $>$ $8.6\times 10^{4}$) compared to raw LaCAM’s solutions.
In summary, layered LaCAM shows little improvement compared to raw LaCAM and has worse solutions due to the introduction of wait actions when merging solutions. However, layered LaCAM has an advantage in memory usage, as it reduces the number of intermediate states by reducing the number of agents solved simultaneously.
VI-H Summary
@@
Moreover, serial MAPF methods (e.g., EECBS, LNS2) generally demand more time compared to parallel MAPF methods (e.g., LaCAM and PIBT+), yet they yield higher-quality solutions.. Notably, serial MAPF methods frequently fail due to time constraints, while parallel MAPF methods tend to fail due to memory limitations.
The observed discrepancy can be attributed to the differing search algorithms utilized: serial methods repeatedly search for complete paths. However, they employ memory-efficient data structures such as priority trees or conflict trees to store knowledge about resolving conflicts. In contrast, parallel methods store previous conflict-solving knowledge as intermediate states to avoid searching for the complete path of agents repeatedly. However, they may exhaust memory resources under complex MAPF instances.
VII Conclusion
Motivated by the exponential growth in the cost of solving MAPF instances (in terms of time and memory usage) as the number of agents increases, we proposed layered MAPF as a solution to reduce the computational burden. This approach decomposes a MAPF instance into multiple smaller subproblems without compromising solvability. Each subproblem is solved in isolation, with consideration given to other subproblems’ solutions as dynamic obstacles. Our methodology involves a progressive decomposition of MAPF instances, ensuring that each step preserves solvability.
In the results of our decomposition of MAPF instances (Section V), we observed that our method is highly effective for MAPF instances with free grids exceeding twice the number of agents. On average, the time cost is around 1s and never exceeds 3s, even for dense instances with 800 to 1000 agents. Memory usage remains below 1MB, with fewer computations and memory space required for maps with more free grids than agents.
When applied to the state-of-the-art methods (EECBS, PBS, LNS, HCA*, Push and Swap, PIBT+, LaCAM), layered MAPF significantly reduces memory usage and time cost, particularly for serial MAPF methods. Consequently, layered MAPF methods achieve higher success rates than raw MAPF methods, especially for serial MAPF. And the quality of solution for the layered version of serial MAPF methods is similar to the raw version, while the layered version of parallel MAPF methods produces inferior solutions due to the introduction of numerous wait actions during solution merging.
@@ VII Conclusion
In the future, we plan to propose new merging solution techniques for parallel methods without compromising solution quality. Additionally, we aim to generalize the idea of decomposing MAPF instances to address extensions of MAPF problems, such as considering the shape of agents.
Acknowledgments
The first author thanks Lu Zhu for her encouragement and support during the consummation of Layered MAPF.
[…]

### v2

@@ 6.4 Push and Swap
Hierarchical Cooperative A* (HCA*) [32] is a suboptimal and incomplete prioritized MAPF algorithm that decouples into a series of single-agent searches. HCA* calculates a distance table for each agent to determine their priority. The distance table stores the distance from every location on the map to the target state. Since HCA* searches one path at a time, and after reviewing its code, we found that there are no other variables consuming significant memory space aside from the distance table, we conclude that HCA*’s memory usage is primarily determined by the size of these distance tables. Specifically, it depends on the number of agents multiplied by the size of the map. Thus, we use the product of the number of agents and the size of the map as an estimate for HCA*’s memory usage.
@@
The average time cost, success rate, memory usage, makespan, and sum of costs for Layered HCA* and raw HCA* are shown in Fig. 23. A comparison of how Layered HCA* and raw HCA* perform under various maps as the number of agents increases is shown in Fig. A5 in the Appendix. As illustrated in Fig. A5, the time cost of Layered HCA* is very close to that of raw HCA*, as HCA* searches each agent’s path separately. However, Layered HCA* has lower memory usage compared to raw HCA*, as it does not need to store the heuristic table for all agents.
In terms of success rate, Layered HCA* has a slightly higher success rate, as HCA* is incomplete, and solving the instance separately may increase its chances of finding a solution (e.g., instances are decomposed into subproblems that involve only one agent).
Regarding path quality, the solution of Layered HCA* is slightly worse than that of raw HCA*. A summary of the performance of Layered HCA* and raw HCA* can be found in Fig. 23.
On average, Layered HCA* has a slightly higher time cost than raw HCA* ($1.06\times 10^{4}$ ms $>$ $1.05\times 10^{4}$ ms) and a similar success rate (0.71 $>$ 0.66) as HCA* solves each agent separately, and decomposition does not improve the efficiency of HCA*. However, Layered HCA* requires less memory space than raw HCA* ($9.1\times 10^{5}$ $<$ $2.6\times 10^{7}$), as Layered HCA* only stores the distance table of the current subproblem’s agents, while raw HCA* needs to store the distance table of all agents.
Regarding path quality, Layered HCA* has a slightly larger makespan ($2.5\times 10^{2}$ $>$ $2.2\times 10^{2}$) and SOC ($5.2\times 10^{4}$ $>$ $4.2\times 10^{4}$). Because HCA* solves agents in order of decreasing distance to the target, raw HCA* solves agents with longer paths first, limiting the makespan of these agents’ solutions, while Layered HCA* breaks this order as agents with long paths may not be in the same subproblem and may not be solved first.
@@
6.6 Priority Inheritance with Backtracking$+$
Priority Inheritance with Backtracking$+$ (PIBT$+$) [16] is an incomplete and suboptimal MAPF algorithm. PIBT$+$ avoids the occurrence of the same state of agents twice in the solution. And the code [^9] https://github.com/Kei18/pibt2 for PIBT$+$ provides no interfaces to avoid conflicts with external paths, making it a parallel MAPF method. In the implementation of PIBT$+$, a function called $funcPIBT$ is used recursively to move all agents one step at a time without any collisions. The function is called repeatedly until all agents reach their targets. In computer systems (e.g., Windows or Linux), function recursion relies on the stack to manage the context of function calls. Each recursive call allocates a new stack frame, and the depth of recursion is limited by the size of the stack. Therefore, the memory usage of PIBT$+$ is primarily determined by the recursion depth of $funcPIBT$. We use the maximum recursion depth as an estimate of the memory usage for PIBT$+$.
The average time cost, success rate, memory usage, makespan, and sum of costs for Layered PIBT$+$ and raw PIBT$+$ are shown in Fig. 23. The performance of Layered PIBT$+$ and raw PIBT$+$ under various maps as the number of agents increases is presented in Fig. A6 in the Appendix.
As shown in Fig. A6, in some cases, Layered PIBT$+$ takes more time compared to raw PIBT$+$, while in other cases, it does not. This is because the time cost of PIBT$+$ is closely tied to the time cost of instance decomposition. Both Layered PIBT$+$ and raw PIBT$+$ have a high success rate, but neither achieves a 100% success rate, as PIBT$+$ is an incomplete MAPF method. In terms of memory usage, Layered PIBT$+$ consistently requires less memory than raw PIBT$+$, even as the number of agents increases.
Regarding path quality, the solutions found by Layered PIBT$+$ are generally worse than those found by raw PIBT$+$, though in some cases, both Layered PIBT$+$ and raw PIBT$+$ solutions see a decrease in SOC and makespan as the number of agents increases. This is because, with more agents, there are fewer opportunities for PIBT$+$ to find longer paths. A summary of Layered PIBT$+$ and raw PIBT$+$ performance can be found in Fig. 23.
In general, Layered PIBT$+$ takes more time than raw PIBT$+$ ($2.2\times 10^{3}$ ms $>$ $1.9\times 10^{3}$ ms). As mentioned in Section 5, the decomposition of MAPF instances with 800 to 1000 agents may cost 1s to 3s, and these instances play a major role in contributing to the mean time cost of the MAPF method. Although decomposition reduces the time cost of PIBT$+$, the time cost of decomposition offsets the reduced time cost and thus increases the overall time cost.
Layered PIBT$+$ has a similar success rate compared to raw PIBT$+$ (Layered PIBT$+$: 0.98, raw PIBT$+$: 0.98), as both Layered PIBT$+$ and raw PIBT$+$ typically take less than 30s in most cases. In terms of memory usage, Layered PIBT$+$ requires less memory space than raw PIBT$+$ ($6.5\times 10^{5}$ $<$ $1.1\times 10^{6}$), as decomposition reduces the number of intermediate states by reducing the number of agents solved simultaneously.
Regarding solution quality, as PIBT$+$ is a parallel MAPF method, Layered PIBT$+$ introduces extra wait actions, resulting in solutions with higher makespan ($9.4\times 10^{3}$ $>$ $4.3\times 10^{2}$) and SOC ($3.2\times 10^{6}$ $>$ $1.3\times 10^{5}$) compared to raw PIBT$+$’s solutions.
@@ 6.6 Priority Inheritance with Backtracking$+$
6.7 Lazy Constraints Addition for MAPF 2
Lazy Constraints Addition for MAPF 2 (LaCAM2) [18] is a complete and suboptimal method. More details about LaCAM2 can be found in Section 2. LaCAM avoids the occurrence of the same state of agents twice in the solution. Compared to LaCAM, LaCAM2 [18] introduces a swap action to the successor generator, allowing for the quick generation of solutions. The code [^10] https://github.com/Kei18/lacam2 for LaCAM2 provides no interfaces to avoid conflicts with external paths, making it a parallel MAPF method. In the implementation, LaCAM2 stores all high-level nodes during planning, which increases rapidly as the number of agents grows. Therefore, LaCAM2’s memory usage is primarily determined by the maximum number of high-level nodes. We use this maximum number as an estimation of LaCAM2’s memory usage.
The average time cost, success rate, memory usage, makespan, and sum of costs for both Layered LaCAM2 and raw LaCAM2 are shown in Fig. 23. The performance of Layered LaCAM2 and raw LaCAM2 across various maps, as the number of agents increases, is presented in Fig. A7 in the Appendix.
As shown in Fig. A7, in some cases, Layered LaCAM2 takes more time than raw LaCAM2, while in other cases it does not, as LaCAM2’s time cost is close to the decomposition time cost of the instance. Both Layered LaCAM2 and raw LaCAM2 achieve a 100% success rate, as LaCAM2 is a complete and highly efficient MAPF method.
In terms of memory usage, Layered LaCAM2 consistently requires less memory than raw LaCAM2, even as the number of agents increases.
Regarding path quality, the solutions generated by Layered LaCAM2 are slightly worse than those generated by raw LaCAM2. A summary of the performance of both Layered LaCAM2 and raw LaCAM2 can be found in Fig. 23.
In average, Layered LaCAM2 has larger time cost than raw LaCAM2 ($9.6\times 10^{2}$ ms $>$ $7.3\times 10^{2}$ ms). Similar to PIBT$+$, the decomposition of MAPF instances with 800 to 1000 agents may cost 1s to 3s, and these instances play a major role in determining the time cost of the MAPF method. Although decomposition reduces the time cost of LaCAM2, the time cost of decomposition offsets the reduced time cost, resulting in no significant difference in the overall time cost.
Both Layered LaCAM2 and raw LaCAM2 achieve the same success rate (Layered LaCAM2: 1.0, raw LaCAM2: 1.0), indicating that both methods successfully find all solutions within 30 seconds. This suggests that the decomposition of the instance does not sacrifice its solvability during the experiments. If there had been any loss of solvability, Layered LaCAM2’s success rate would have been lower than 1.0. Based on this data, we speculate that the likelihood of our decomposition causing a loss of solvability is less than 1%. In terms of memory usage, Layered LaCAM2 requires significantly less memory space than raw LaCAM2 ($3.4\times 10^{4}$ configs $<$ $1.7\times 10^{5}$ configs), as decomposition reduces the number of intermediate states by reducing the number of agents solved simultaneously.
Regarding solution quality, Layered LaCAM2 introduces extra wait actions, resulting in solutions with higher makespan ($9.5\times 10^{3}$ $>$ $3.4\times 10^{2}$) and SOC ($3.2\times 10^{6}$ $>$ $9.5\times 10^{4}$) compared to raw LaCAM2’s solutions.
LaCAM2 is the only algorithm among the seven that achieves a 100% success rate, making it a useful benchmark for estimating the loss of solvability caused by our decomposition of the instance. Since Layered LaCAM2 also maintains a 100% success rate, we speculate that the loss of solvability due to our decomposition is less than 1%.
In summary, Layered LaCAM2 shows little improvement compared to raw LaCAM2 and has worse solutions due to the introduction of wait actions when merging solutions. However, Layered LaCAM2 has an advantage in memory usage, as it reduces the number of intermediate states by reducing the number of agents solved simultaneously.
6.8 Summary
@@
Moreover, serial MAPF methods (e.g., EECBS, LNS2) generally demand more time compared to parallel MAPF methods (e.g., LaCAM and PIBT+), yet they yield higher-quality solutions. Notably, serial MAPF methods frequently fail due to time constraints, while parallel MAPF methods tend to fail due to memory limitations.
The observed discrepancy can be attributed to the differing search algorithms utilized: serial methods repeatedly search for complete paths. However, they employ memory-efficient data structures such as priority trees or constraint trees to store knowledge about resolving conflicts. In contrast, parallel methods store previous conflict-solving knowledge as intermediate states to avoid searching for the complete path of agents repeatedly. However, they may exhaust memory resources under complex MAPF instances.
7 Conclusion
MAPF is extensively utilized in autonomous systems, such as automated warehouses, UAV traffic management and multi-agent exploration. Motivated by the exponential growth in the cost of solving MAPF instances (in terms of time and memory usage) as the number of agents increases, we proposed Layered MAPF as a solution to reduce the computational burden. This approach decomposes a MAPF instance into multiple smaller subproblemsprogressively, and minimize the possibility of loss of solvability by introducing legality check during each step of decomposition. In terms of time complexity, in the worst case, time complexity of decomposition is $\mathcal{O}(k^{5})$; while in the best case, its time complexity is $\mathcal{O}(k^{3})$, where $k$ represents the number of agents.
In the results of our decomposition of MAPF instances (Section 5), we observed that our method is highly effective for MAPF instances with passable cells exceeding twice the number of agents. On average, the time cost is around 1s and never exceeds 3s, even for dense instances with 800 to 1000 agents. Memory usage remains below 1MB, with fewer computations and memory space required for maps with more passable cells than agents.
When applied to the state-of-the-art methods (EECBS, PBS, LNS2, HCA*, Push and Swap, PIBT+, LaCAM2), Layered MAPF significantly reduces memory usage and time cost, particularly for serial MAPF methods. Consequently, Layered MAPF methods achieve higher success rates than raw MAPF methods, especially for serial MAPF. And the quality of solution for the Layered version of serial MAPF methods is similar to the raw version, while the Layered version of parallel MAPF methods produces inferior solutions due to the introduction of numerous wait actions during solution merging.
@@ 7 Conclusion
In the future, we plan to propose new merging solution techniques for parallel methods without compromising solution quality. Essentially, the decomposition of the instance is a trade-off between solution quality and efficiency, and a systematic exploration of this trade-off between runtime and solution quality is necessary. Given that the initial state of the cluster plays a crucial role in the decomposition process, a randomized restart approach could benefit the decomposition. We plan to generate different initial clusters during the bipartition phase and select the one that results in the best decomposition. While this process introduces some extra time cost, it may lead to better decomposition outcomes. Additionally, we aim to generalize the idea of decomposing MAPF instances to address extensions of MAPF problems, such as considering the shape of agents.
Appendix A Comparison between Layered MAPF and raw MAPF as the number of agent increasing
Figure A1: These figures illustrate how Layered EECBS and raw EECBS work under various maps as the number of agents increases. They are compared in terms of time cost, success rate, memory usage, and solution quality (sum of cost and makespan). The data for Layered EECBS are shown with a solid line, and the data for raw EECBS are shown with a dotted line. More details about the maps can be found in Fig. 15 and Fig. 16.
Figure A2: These figures illustrate how Layered PBS and raw PBS work under various maps as the number of agents increases. They are compared in terms of time cost, success rate, memory usage, and solution quality (sum of cost and makespan). The data for Layered PBS are shown with a solid line, and the data for raw PBS are shown with a dotted line. More details about the maps can be found in Fig. 15 and Fig. 16.
Figure A3: These figures illustrate how Layered LNS2 and raw LNS2 works under various of maps as number of agents increasing, they are compared in term of time cost, success rate, memory usage and solution quality (sum of cost and makespan). Layered LNS2’s data are shown in solid line and raw LNS2’s data are shown in dotted line. More details about maps can be found in Fig.15 and Fig. 16.
Figure A4: These figures illustrate how Layered Push and Swap and raw Push and Swap work under various maps as the number of agents increases. They are compared in terms of time cost, success rate, memory usage, and solution quality (sum of cost and makespan). The data for Layered Push and Swap are shown with a solid line, and the data for raw Push and Swap are shown with a dotted line. More details about the maps can be found in Fig. 15 and Fig. 16.
Figure A5: These figures illustrate how Layered HCA and raw HCA work under various maps as the number of agents increases. They are compared in terms of time cost, success rate, memory usage, and solution quality (sum of cost and makespan). The data for Layered HCA are shown with a solid line, and the data for raw HCA are shown with a dotted line. More details about the maps can be found in Fig. 15 and Fig. 16.
Figure A6: These figures illustrate how Layered PIBT+ and raw PIBT+ work under various maps as the number of agents increases. They are compared in terms of time cost, success rate, memory usage, and solution quality (sum of cost and makespan). The data for Layered PIBT+ are shown with a solid line, and the data for raw PIBT+ are shown with a dotted line. More details about the maps can be found in Fig. 15 and Fig. 16.
Figure A7: These figures illustrate how Layered LaCAM2 and raw LaCAM2 work under various maps as the number of agents increases. They are compared in terms of time cost, success rate, memory usage, and solution quality (sum of cost and makespan). The data for Layered LaCAM2 are shown with a solid line, and the data for raw LaCAM2 are shown with a dotted line. More details about the maps can be found in Fig. 15 and Fig. 16.